### v2

@@ Abstract
Unsupervised visual clustering has garnered significant attention in recent times, aiming to characterize distributions of unlabeled visual images through clustering based on a parameterized appearance approach. Alternatively, clustering algorithms can be viewed as assignment problems, often characterized as NP-hard, yet precisely solvable for small instances on contemporary hardware. Adiabatic quantum computing (AQC) emerges as a promising solution, poised to deliver substantial speedups for a range of NP-hard optimization problems. However, existing clustering formulations face challenges in quantum computing adoption due to scalability issues. In this study, we present the first clustering formulation tailored for resolution using Adiabatic quantum computing. An Ising model is introduced to represent the quantum mechanical system implemented on AQC. The proposed approach demonstrates high competitiveness compared to state-of-the-art optimization-based methods, even when utilizing off-the-shelf integer programming solvers. Lastly, this work showcases the solvability of the proposed clustering problem on current-generation real quantum computers for small examples and analyzes the properties of the obtained solutions.
@@
clustering, quantum computing, quantum annealing, neural network
@@ 1 Introduction
Unsupervised learning in automated object and human understanding has recently become one of the most popular research topics in robot vision. It is primarily due to the abundance of unlabelled raw data and the need for robust visual recognition algorithms that can perform consistently in various challenging conditions. However, traditional methods, like K-Mean clustering, have shown limitations regarding speed and accuracy when dealing with large-scale databases. These methods heavily rely on a predetermined number of clusters ($k$), which is not applicable in real scenarios, such as a visual object or visual landmark recognition clustering. In recent years, several studies [1, 2, 3] based on deep learning have been proposed to address the challenge of unsupervised clustering. However, these methods still rely on rule-based algorithms for the final cluster construction, and since they need to process large-scale databases, this step still needs to be completed.
In this paper, we will map the clustering problem into a quantum mechanical system whose energy is equivalent to the cost of the optimization problem. Therefore, if it is possible to measure the lowest energy state of the system, a solution to the corresponding optimization problem is found. It is done with an Adiabatic Quantum Computer (AQC), which implements a quantum mechanical system made from qubits and can be described by the Ising model. Using this approach, a quantum speedup, which further scales with system size and temperature, has already been shown for applications in physics.
While quantum computing can provide a range of future advantages, mapping a problem to an AQC is nontrivial. It often requires reformulating the problem from scratch, even for well-investigated tasks. On the one hand, the problem needs to be matched to the Ising model. On the other hand, real quantum computers have a minimal number of qubits and are still prone to noise, which requires tuning the model to handle the limitations. In this work, we present the first quantum computing approach for visual clustering.
@@
  • Initially, we formulate the clustering problem as a Quadratic Unconstrained Binary Optimization (QUBO) problem. Different from previous methods, our formulation can work even number of clusters is unspecified.
  • Lastly, we present an efficient solution approach for this QUBO problem using Adiabatic Quantum Computing, aiming to attain promising unsupervised clustering outcomes.
Throughout the rest of this paper, we will present related studies addressing the visual clustering problem. Subsequently, we articulate the visual clustering problem within the quantum framework. This will be followed by an exploration of fundamental concepts in quantum computing. We will then detail our approach to framing the visual clustering problem through a QUBO formulation. Lastly, we demonstrate the outcomes of our experiments conducted on the D-Wave quantum computer and draw comparisons with the $k$-NN clustering algorithm.
@@ 2 Related Work
Quantum Machine Learning: Quantum machine learning algorithms have been recently developed to solve problems that are too slow to solve on classical computers. Examples include Shor’s Algorithm to factor large integers [4], Grover’s Algorithm for searching [5], and the Deutsch-Josza algorithm for determining whether a black box function is constant or balanced [6].
A class of discrete optimization problems known as Quadratic Unconstrained Binary Optimization problems is particularly amenable to a technique known as Adiabatic Quantum computing. This paper investigates how classical computing and Adiabatic Quantum computing can be unified to solve the Visual Clustering problem by formulating it in an end-to-end QUBO framework. Prior studies have used the strategy of solving QUBO problems with Adiabatic Quantum computing for tasks like cluster assignment, robust fitting, and $k$-means clustering. [7, 8, 9]
Deep Visual Clustering: Visual clustering is a crucial tool for data analytics. Deep visual clustering refers to applying deep learning methods to clustering. Classical techniques for clustering such as density-based clustering, centroid-based clustering, distribution-based clustering, ensemble clustering, etc. [10] display limited performance on complex datasets. Deep neural networks seek to map complex data to a feature space where it becomes easier to cluster. [10] Modern deep clustering methods have focused on Graph Convolutional Neural Networks (GCN) and Transformers [11, 12, 13].
Empirical data is often most naturally represented with a graph structure. For example, websites are usually connected by a hyperlink. Social media platform users are connected by being followers or stores connected by roads. GCNs exploit this graph structure by taking it as input to produce more accurate features, which are used finally to cluster the nodes. [14, 15, 16, 17] Transformers also exploit the connections between the data; Ling et al. unified for the first time the transformer model and contrastive learning for image clustering. [13] Nguyen et al. presented the Clusformer model, which addresses the sensitivity of the GCN to noise and uses a transformer for the task of visual clustering [18].
@@ 3.2 Quantum Superposition
$\begin{split}q_{1}&=\ket{0}\\ q_{2}&=\frac{1}{\sqrt{2}}\ket{0}+\frac{1}{\sqrt{2}}\ket{1}\end{split}$
@@ 3.5 Adiabatic Quantum Computing
The quantum state |$\ket{\psi(t)}$ of a system of qubits varies with time according to the Schrödinger equation:
@@
where $H(t)$ is an operator called the Hamiltonian of the quantum system. At each time $t$, $H(t)$ describes the energy of the quantum state. Encoding the QUBO problem in a Hamiltonian $H_{QUBO}(t)$ means that the answer is the state of the qubits when $H_{QUBO}(t)$ is in the lowest energy state. Initializing $H_{QUBO}$ in its lowest energy state is problematic because it is complex. Many researchers use the strategy of Adiabatic Quantum Computing to compute the lowest energy state of a complex quantum system, which is where we initialize the quantum system in the lowest energy state of some easy-to-compute Hamiltonian $H_{BASIC}$, and form our Hamiltonian as the interpolation between them:
@@
4 Deep Visual Clustering
Let $\mathcal{D}$ be a set of $N$ data points to be clustered. We define $\mathcal{H}:\mathcal{D}\rightarrow\mathcal{X}\subset\mathbb{R}^{N\times d}$ as a function that embeds these data points to a latent space $\mathcal{X}$ of $d$ dimensions. Let $\hat{Y}$ be a set of ground-truth cluster IDs assigned to the corresponding data points. Since $\hat{Y}$ has a maximum of $N$ different cluster IDs, we can represent $\hat{Y}$ as a vector of $\left[0,N-1\right]^{N}$. A deep clustering algorithm $\Phi:\mathcal{X}\rightarrow Y\in\left[0,N-1\right]^{N}$ is defined as a mapping that maximizes a clustering metric $\sigma(Y,\hat{Y})$.
As $\hat{Y}$ is not accessible during the training stage, the number of clusters and samples per cluster are not given beforehand. Therefore, divide and conquer philosophy can be adapted to relax the problem. In particular, a deep clustering algorithm $\Phi$ can be divided into three sub-functions: pre-processing $\mathcal{K}$, in-processing $\mathcal{M}$, and post-processing $\mathcal{P}$ functions. Formally, $\Phi$ can be presented as in Eqn. (1).
$\Phi(\mathcal{X})=\{\mathcal{M}\circ\mathcal{K}(\mathbf{x}_{i},n)\}_{\mathbf{x}_{i}\in\mathcal{X}}$ (1)
where $\mathcal{K}$ denotes an unsupervised clustering algorithm on deep features $\mathbf{x}_{i}$ of the $i$-th sample. k-NN is usually a common choice for $\mathcal{K}$ [2, 1, 21, 22, 3].
In addition, for the post-processing $\mathcal{P}$ function, it is usually a rule-based algorithm to merge two or more clusters into one when they share high concavities [1, 2, 21]. Since $\mathcal{P}$ does not depend on the input data points nor is a learnable function, the parameters $\theta_{\mathcal{M}}$ of $\Phi$ can be optimized via this objective function:
$\theta^{*}_{\mathcal{M}}=\arg\min_{\theta_{\mathcal{M}}}\mathbb{E}_{\mathbf{x}_{i}\sim p(\mathcal{X})}\left[\mathcal{L}(\mathcal{M}\circ\mathcal{K}(\mathbf{x}_{i},n),\hat{q}_{i})\right]$ (2)
where $k$ is the number of the nearest neighbors of input $\mathbf{x}_{i}$, $\hat{q_{i}}$ is the objective of $\mathcal{M}$ to be optimized, and $\mathcal{L}$ denotes a suitable clustering or classification loss.
@@ 4 Deep Visual Clustering
Given that the clustering method typically consists of three components, namely $\mathcal{K}$, $\mathcal{M}$, and $\mathcal{P}$, previous studies primarily concentrated on enhancing the performance of $\mathcal{M}$. However, the $\mathcal{P}$ also holds an important role contributing to the overall performance [18]. In most studies [18, 23, 24], the design of $\mathcal{P}$ is typically addressed with a simple rule-based algorithm, which often proves to be sensitive and reliant on the threshold chosen. Consequently, this sensitivity significantly impacts the overall performance of the clustering system, rendering it unstable. In this paper, we thoroughly investigate the design of $\mathcal{P}$ from an optimization perspective. To address this challenge, we propose a novel approach that leverages quantum machine techniques, improving unsupervised clustering performance.
Figure 1: A visualization of the problem formulation that shows the flow of images from the image set to their final assignment to an anchor cluster.
5 Our Proposed Method
5.1 QUBO Formulation
@@
In this section, we formulate the $\mathcal{P}$ component from the optimization perspective. Given a set of clusters $\mathcal{G}=\{g_{0},g_{1},\dots,g_{u-1}\}$. It is noted that $g_{i}=\mathcal{M}(\mathcal{K}(\mathbf{x}_{i},n)$) is processed by deep network $\mathcal{M}$ to remove the noisy samples, thus, number of samples in each cluster $|g_{i}|$ can vary among clusters. The objective of $\mathcal{P}$ can be defined as in Eqn. (3).
@@
where $\mathbbm{1}$ is the indicator function, $c_{i}$ is the cluster id of cluster $g_{i}$. Typically, the $c_{i}$ can be determined by the centroid $\mathbf{x}_{i}$. This information is available in the training phase, but in the testing phase. For this reason, $\mathcal{P}$ has to determine which clusters share the same ID. To address this problem, we find the optimal solution of Eqn. (3) using the Quadratic Unconstrained Binary Optimization Problem.
We can divide the $\mathcal{G}$ into two sets of clusters, denoted as $\mathcal{A}$ and $\mathcal{B}$. Let $\mathcal{A}=\{\textbf{a}_{1},\textbf{a}_{2},\dots\textbf{a}_{r}\}$ be a set of anchor clusters where each individual cluster $\textbf{a}_{i}$ is assigned a distinct ID. Let =$\mathcal{B}=\{\textbf{b}_{1},\textbf{b}_{2},\dots\textbf{b}_{s}\}$ be the set of unknown clusters where we do not know their IDs. Our goal is to assign $\textbf{b}_{j}$ to a cluster $\textbf{a}_{i}$ if they share the same ID or add $\textbf{b}_{j}$ to be a new element of $\mathcal{A}$ if it does not match with any cluster inside $\mathcal{A}$. Formally, let $\mathbf{u}\in\mathbbm{R}^{r\times s}$ be the binary matrix where,
@@
Since $\textbf{b}_{j}$ could not match more than one $\textbf{a}_{i}$. Thus we have a constraint as presented in Eqn. (4).

### v3

@@ Abstract
Unsupervised visual clustering has garnered significant attention recently, aiming to characterize distributions of unlabeled visual images through clustering based on a parameterized appearance approach. Alternatively, clustering algorithms can be viewed as assignment problems where an unlabelled sample needs to be assigned to a specific cluster. This problem can be formulated as Quadratic Unconstrained Binary Optimization that is characterized as NP-hard, yet precisely solvable for small instances on contemporary hardware. Adiabatic quantum computing (AQC) emerges as a promising solution to offer a good speed in solving the NP-hard optimization problems. However, existing clustering approaches face challenges in designing the problems to be solved by the quantum computer. In this study, we present the first clustering formulation tailored for resolution using Adiabatic quantum computing. An Ising model is introduced to represent the quantum mechanical system implemented on AQC. The proposed approach demonstrates high competitiveness compared to state-of-the-art optimization-based methods, even when utilizing off-the-shelf integer programming solvers. Lastly, this work showcases the solvability of the proposed clustering problem on current-generation real quantum computers for small examples and analyzes the properties of the obtained solutions.
@@
Quantum Computing, Visual Clustering, Quantum Annealing, Neural Networks
Article Highlights
  • The paper presents a novel Quadratic Unconstrained Binary Optimization (QUBO) formulation for unsupervised clustering.
  • The paper leverages quantum mechanics to solve the QUBO problem.
  • The empirical experiments demonstrate the efficientness of the quantum computers in solving unsupervised clustering problems.
@@ 1 Introduction
Unsupervised learning in automated object and human understanding has recently become one of the most popular research topics in robot vision. It is primarily due to the abundance of unlabelled raw data and the need for robust visual recognition algorithms that can perform consistently in various challenging conditions. Traditional methods, like K-Mean clustering, have shown limitations regarding speed and accuracy when dealing with large-scale databases. These methods heavily rely on a predetermined number of clusters ($k$), which is not applicable in real scenarios, such as a visual object or visual landmark recognition clustering. In recent years, several studies [1, 2, 3] based on deep learning have been proposed to address the challenge of unsupervised clustering. However, these methods rely on rule-based algorithms for the final cluster construction. Meaning they are still required to take additional steps to process large-scale databases.
In this paper, we will map the clustering problem into a quantum mechanical system whose energy is equivalent to the cost of the optimization problem. Therefore, if it is possible to measure the lowest energy state of the system, a solution to the corresponding optimization problem can be derived. It is done with an Adiabatic Quantum Computer (AQC), which implements a quantum mechanical system made from qubits and can be described by the Ising model. Using this approach, a quantum speedup, which further scales with system size and temperature, has already been shown for applications in physics.
While quantum computing can provide a range of future advantages, formulating a problem in the form of AQC is a nontrivial task. It often requires reformulating the problem from scratch, even for well-investigated tasks. Since the problem needs to be matched to the Ising model, real quantum computers have a minimal number of qubits and are still prone to noise, which requires tuning the model to handle the limitations. In this work, we present the first quantum computing approach for visual clustering.
@@
  • Initially, we formulate the clustering problem as a Quadratic Unconstrained Binary Optimization (QUBO) problem. Unlike previous methods, our formulation allows us not to need to know the number of clusters.
  • We present an efficient solution approach for the QUBO problem using Adiabatic Quantum Computing, aiming to attain promising unsupervised clustering outcomes.
  • The results from the experiment illustrate the feasibility of our proposed solution for the clustering problem.
Throughout the rest of this paper, we will present related studies addressing the visual clustering problem. Subsequently, we articulate the visual clustering problem within the quantum framework. It will be followed by an exploration of fundamental concepts in quantum computing. We will then detail our approach to framing the visual clustering problem through a QUBO formulation. Lastly, we demonstrate the outcomes of our experiments conducted on the D-Wave quantum computer and draw comparisons with the $k$-NN clustering algorithm.
@@ 2 Related Work
Quantum Machine Learning: Quantum machine learning algorithms have been recently developed to solve problems that are too slow to solve on classical computers [4, 5, 6, 7] A class of discrete optimization problems known as Quadratic Unconstrained Binary Optimization problems is particularly amenable to a technique known as Adiabatic Quantum computing. This paper investigates how classical computing and Adiabatic Quantum computing can be unified to solve the Visual Clustering problem by formulating it in an end-to-end QUBO framework. Prior studies have used the strategy of solving QUBO problems with Adiabatic Quantum computing for tasks like cluster assignment, robust fitting, and $k$-means clustering. [8, 9, 10]
Deep Visual Clustering: Visual clustering is a crucial tool for data analytics. Deep visual clustering refers to applying deep learning methods to clustering. Classical techniques for clustering such as density-based clustering, centroid-based clustering, distribution-based clustering, ensemble clustering, etc. [11] display limited performance on complex datasets. Deep neural networks seek to map complex data to a feature space where it becomes easier to cluster. [11] Modern deep clustering methods have focused on Graph Convolutional Neural Networks (GCN) and Transformers [12, 13, 14]. Empirical data is often most naturally represented with a graph structure. For example, websites are usually connected by a hyperlink. Social media platform users are connected by being followers or stores connected by roads. GCNs exploit this graph structure by taking it as input to produce more accurate features, which are used finally to cluster the nodes. [15, 16, 17, 18] Transformers also exploit the connections between the data; Ling et al. unified for the first time the transformer model and contrastive learning for image clustering. [14] Nguyen et al. presented the Clusformer model, which addresses the sensitivity of the GCN to noise and uses a transformer for the task of visual clustering [19].
@@ 3.2 Quantum Superposition
$\begin{split}q_{1}&=\frac{1}{\sqrt{2}}\ket{0}+\frac{1}{\sqrt{2}}\ket{1}\\ q_{2}&=\ket{0}\end{split}$
@@ 3.5 Adiabatic Quantum Computing
In the previous sections, we introduced fundamental quantum primaries. In this section, we present how quantum annealing works in D-wave. The quantum state |$\ket{\psi(t)}$ of a system of qubits varies with time according to the Schrödinger equation:
@@
where $H(t)$ is an operator called the Hamiltonian of the quantum system. At each time $t$, $H(t)$ describes the energy of the quantum state. Encoding the QUBO problem in a Hamiltonian $H_{QUBO}(t)$ means that the answer is the state of the qubits when $H_{QUBO}(t)$ is in the lowest energy state. Since we do not know the lowest energy state of the QUBO prior, we could not initialize the $H_{QUBO}$ in that state. Many researchers use the strategy of Adiabatic Quantum Computing to compute the lowest energy state of a complex quantum system, which is where we initialize the quantum system in the lowest energy state of some easy-to-compute Hamiltonian $H_{BASIC}$, and form our Hamiltonian as the interpolation between them:
@@
In the D-wave quantum computer, these processes are produced by the following ideas. First, the system begins with qubits, each in a superposition of 0 and 1, initially uncoupled. During quantum annealing, couplers and qubit biases are introduced, leading to entanglement among the qubits. This entangled state represents a multitude of possible solutions. By the end of the annealing process, each qubit settles into a classical state corresponding to the problem’s minimum energy configuration or something very close. In D-Wave quantum computers, this entire process occurs within microseconds.
4 Problem Definition and Motivations
4.1 Deep Visual Clustering
Let $\mathcal{D}$ be a set of $N$ data points to be clustered. We define an embedding function $\mathcal{H}:\mathcal{D}\to\mathbb{R}^{d}$ that maps each data point $x_{i}\in\mathcal{D}$ to a latent space $\mathcal{X}\subset\mathbb{R}^{d}$. That is, each sample is transformed as:
$\mathbf{x}_{i}=\mathcal{H}(x_{i}),\quad\forall x_{i}\in\mathcal{D}.$ (1)
Let $\hat{y}_{i}$ be the ground-truth cluster ID assigned to the data point $x_{i}$. Since there are at most $N$ unique clusters, each $\hat{y}_{i}$ is an integer in the range $[0,N-1]$, forming a vector $\hat{Y}=(\hat{y}_{1},\hat{y}_{2},...,\hat{y}_{N})\in[0,N-1]^{N}$. A deep clustering algorithm $\Phi$ is then defined as a function that assigns a cluster label $y_{i}$ to each embedded data point $\mathbf{x}_{i}$, maximizing a clustering metric $\sigma(y_{i},\hat{y}_{i})$. Since the ground-truth labels $\hat{y}_{i}$ are inaccessible during training, the number of clusters and samples per cluster are unknown beforehand. To address this challenge, a divide-and-conquer strategy is employed, decomposing $\Phi$ into three sub-functions: Pre-processing $\mathcal{K}$: Applies unsupervised clustering at a local level. In-processing $\mathcal{M}$: Learns feature-based clustering representations. Post-processing $\mathcal{P}$: Merges clusters based on structural properties.
Thus, for each data point $\mathbf{x}_{i}$, the clustering function $\Phi$ is defined as:
$\Phi(\mathbf{x}_{i})=\mathcal{P}(\mathcal{M}(\mathcal{K}(\mathbf{x}_{i},k))),$ (2)
where $\mathcal{K}(\mathbf{x}_{i},k)$ retrieves $k$ nearest neighbors to form an initial local grouping. $\mathcal{M}$ refines the grouping through a learnable clustering process, and $\mathcal{P}$ further refines clusters via rule-based merging.
Since $\mathcal{P}$ is typically a heuristic, non-learnable function [1, 2, 22], optimization focuses on $\mathcal{M}$, leading to the objective:
$\theta^{*}_{\mathcal{M}}=\arg\min_{\theta_{\mathcal{M}}}\mathbb{E}_{\mathbf{x}_{i}\sim p(\mathcal{X})}\left[\mathcal{L}(\mathcal{M}(\mathcal{K}(\mathbf{x}_{i},k)),\hat{q}_{i})\right],$ (3)
where $\hat{q}_{i}$ represents the target clustering structure, and $\mathcal{L}$ is a suitable clustering or classification loss function.
4.2 Motivations
@@
Given that the clustering method typically consists of three components, namely $\mathcal{K}$, $\mathcal{M}$, and $\mathcal{P}$, previous studies primarily concentrated on enhancing the performance of $\mathcal{M}$. However, the $\mathcal{P}$ also holds an important role contributing to the overall performance [19]. In most studies [19, 23, 24], the design of $\mathcal{P}$ is typically addressed with a simple rule-based algorithm, which often proves to be sensitive and reliant on the threshold chosen. Consequently, this sensitivity significantly impacts the overall performance of the clustering system, rendering it unstable. In this paper, we thoroughly investigate the design of $\mathcal{P}$ from an optimization perspective. To address this challenge, we propose a novel approach that leverages quantum machine techniques, improving unsupervised clustering performance. The overall architecture is illustrated in Fig 1.
Figure 1: Overall system architecture. The blue box indicates the deep visual feature extraction step. Specifically, $\mathcal{D}$ is the set of N data points. $\mathcal{H}$ is the feature extraction module and $\mathcal{X}$ is the corresponding feature set of the datapoint $\mathcal{D}$. The red box is the deep unsupervised algorithm where $\mathcal{K}$ denotes the K-NN algorithm, $\mathcal{M}$ is the deep cluster algorithm, i.e., Clusformer and $\mathcal{P}$ indicates the post-unsupervised clustering step. The green box indicates our focus in this paper.
5 Our Proposed Method - Quadratic Unconstrained Binary Optimization Problem Approach
@@
In this section, we formulate the $\mathcal{P}$ component from the optimization perspective. Given a set of clusters $\mathcal{G}=\{g_{0},g_{1},\dots,g_{u-1}\}$. It is noted that $g_{i}=\mathcal{M}(\mathcal{K}(\mathbf{x}_{i},k)$) is processed by deep network $\mathcal{M}$ to remove the noisy samples, thus, number of samples in each cluster $|g_{i}|$ can vary among clusters. In short, $g_{i}$ is generated and corresponding to $x_{i}$. In particular, given $x_{i}$, we first apply $\mathcal{K}$ (K-NN algorithm) to obtain $k$ nearest neighbors to form a cluster where $x_{i}$ holds the role as centroid. However, these clusters might have incorrect samples. Therefore, we apply Clusformer (denoted as $\mathcal{M}$) to detect and clear these samples. Finally, we obtain $g_{i}=\mathcal{M}(\mathcal{K}(x_{i},k))$ is a cleaned cluster where $x_{i}$ is a centroid. The objective of $\mathcal{P}$ can be defined as in Eqn. (4).
@@
where $\mathbbm{1}$ is the indicator function, $c_{i}$ is the cluster id of cluster $g_{i}$. Typically, the $c_{i}$ can be determined by the centroid $\mathbf{x}_{i}$. This information is available in the training phase, but not in the testing phase. For this reason, $\mathcal{P}$ has to determine which clusters share the same ID. To address this problem, we find the optimal solution of Eqn. (4) using the Quadratic Unconstrained Binary Optimization Problem (QUBO).
We can divide the $\mathcal{G}$ into two sets of clusters, denoted as $\mathcal{A}$ and $\mathcal{B}$. Let $\mathcal{A}=\{\textbf{a}_{1},\textbf{a}_{2},\dots\textbf{a}_{r}\}$ be a set of anchor clusters where each individual cluster $\textbf{a}_{i}$ is assigned a distinct ID. Let =$\mathcal{B}=\{\textbf{b}_{1},\textbf{b}_{2},\dots\textbf{b}_{s}\}$ be the set of unknown clusters where we do not know their IDs. Our goal is to assign $\textbf{b}_{j}$ to a cluster $\textbf{a}_{i}$ if they share the same ID or add $\textbf{b}_{j}$ to be a new element of $\mathcal{A}$ if it does not match with any cluster inside $\mathcal{A}$. Formally, let $\mathbf{u}\in\mathbbm{R}^{r\times s}$ be the binary matrix where r and s are the number of clusters of $\mathcal{A}$ and $\mathcal{B}$. Each element $_{ij}\in\textbf{u}$ u either 0 or 1. Specially,.
@@
To find the matrix $\mathbf{u}$, the formulation of the unsupervised problem includes following constraints.
5.1 Cluster Constraint
Since $\textbf{b}_{j}$ could not match more than one $\textbf{a}_{i}$, this constraint is represented in Eqn. (5).